# The Streaming Batch Model for Efficient and Fault-Tolerant Heterogeneous Execution

### v4

@@ 3 Overview: The Streaming Batch Model
The streaming batch model (Figure 2(c)) uses task-based execution with a centralized scheduler. Tasks can run on any executor, similar to batch processing (Figure 2(a)). However, tasks across different stages are pipelined and dynamically repartitioned (Figure 3(b)), similar to stream processing systems (Figure 2(b)). Table 1 shows a full feature comparison.
There are two primary challenges: (1) extending lineage reconstruction to support streaming batch execution, and (2) building a centralized scheduler that adaptively manages all running tasks and intermediate partitions.
Challenge 1: Fault-tolerant, memory-efficient partitioning.
Batch processing tasks are often memory-inefficient when operators require different parallelisms. For example, Figure 3(a) shows a two-stage pipeline where parallelism=1 and 3 for stages A and B, respectively. Repartitioning A1 requires a stage barrier, causing high memory usage.
Our goal is to achieve the memory efficiency of stream processing but maintain the low run-time and recovery overheads of lineage-based recovery. In particular, we want to allow the user to set a single target partition size, instead of requiring the user to estimate the number of partitions per operator. This is challenging because then the number of partitions is unknown until run time. Meanwhile, lineage-based systems typically require logging the operations before execution.
To address this, we propose a streaming repartition (§ 4.2.1): a task may dynamically generate multiple output partitions, based on its real-time memory consumption. For example, in Figure 3(b), if A1 experiences memory pressure, it can output a partition early. As we also support pipelined stage execution, this reduces peak memory usage: B tasks can begin and release their input partitions while A1 is executing.
For recovery, we extend Ray’s lineage-based recovery [44] (§ 4.2.2). Ray includes a distributed object store, which we use for intermediate partitions, and task-parallel execution. Like other lineage-based systems, Ray requires immutable task definitions and does not allow task outputs to be read before task completion. To support streaming repartition, we add support in Ray for tasks with dynamic and streaming outputs.
(a) Repartitioning in batch processing.
(b) Streaming repartition.
Figure 3: Streaming repartition (b) allows executors to: (1) locally decide when to output partitions, and (2) pipeline execution with the next operator. Overall, this reduces peak memory usage compared to repartitioning in batch processing systems (a).
@@
Figure 4: Scheduling under memory pressure. Green represents CPU executors’ local memory capacity (1 partition per CPU, 2 total). Pink represents the system’s shared memory capacity for intermediate data (1 partition total). (1): If shared memory is full, executors must stall and buffer outputs locally until space is made, either via spilling to disk or executing downstream tasks. (b) is faster because it schedules A3 as soon as possible but doesn’t stall CPU0.
Challenge 2: Dynamic reconfiguration.
One advantage of the streaming batch model is that it uses a centralized scheduler to schedule all tasks, giving the scheduler a global view of all resources. This enables enforcing a hard limit on memory used by intermediate data partitions. However, this must be done without unnecessarily stalling compute tasks. Memory is shared among all operators, and oversubscribing memory can cause extreme slowdowns from stalls or spilling to disk. When memory is limited, the choice of which task to execute and when is not obvious.
For example, consider Figure 4(a). Each CPU executor has local memory capacity (green) for one partition, and the system has shared memory capacity (pink) for one partition. In phase (1), A2 must stall and buffer its outputs locally until B1 completes. A conservative scheduler additionally waits for phase (2) before scheduling A3, to avoid stalling CPU0. Scheduling A3 optimistically so that it finishes simultaneously with B2, reduces overall run time (Figure 4(b)). Applying such optimizations requires dynamic profiling and reconfiguration.
To address this challenge, we introduce an adaptive scheduler in Section 4.3. The key idea is to fairly allocate shared compute resources between operators, while estimating future memory availability from run-time information to enable optimistic scheduling.
@@
materialize Materialize all items.
@@
3.1 The Dataset API
A Dataset represents an application pipeline. Datasets are lazily created, by reading files or applying transforms to an existing Dataset (Table 2). A Dataset is materialized through a write operation, e.g., to cloud storage, or by iterating over the items in memory.
A key part of the API is the ability to express resource requirements. Resource requirements are a map from resource name to float value and may be passed as an option to the transforms. By default, each transform requires 1 CPU. Resource names can be CPU, GPU, or a custom resource label.
Most of the map-style transforms take a stateless and pure user-defined function (UDF) as an argument. For operations that require significant initialization time, such as a model loaded into GPU memory, we also support stateful UDFs that can be instantiated once and called multiple times on different items. We assume that all UDFs are pure, to enable lineage-based recovery.
Figure 5: Ray Data architecture overview. Ray Data executes as a Ray library. The Ray Data scheduler executes as a Ray driver, dispatching tasks to Ray workers (dashed arrows).
@@ 3.2 Executing a Ray Data Program
After creating a Dataset, the user triggers execution by calling one of the consumption APIs, as seen in (1) in Figure 5. The Dataset is represented as a DAG of logical operators, as in Figure 1. The system’s query planner then compiles this logical DAG into a DAG of physical operators ((2) in Figure 5). The query planner applies operator fusion and decides the number of partitions to use for the first operator (§ 4.1). Each physical operator defines a transform to apply to each row and metadata such as the resources required. Physical operators are decoupled from executors, and their parallelism may not be determined until run time.
Ray Data uses Ray as a task backend, storing intermediate data partitions in Ray’s distributed object store. During execution, the Ray Data scheduler dispatches each physical operator as one or more Ray tasks ((3) in Figure 5). Stateless UDFs are executed as Ray tasks, which are Python processes that can run anywhere in the cluster. Stateful UDFs are executed as tasks running on Ray actors, which are long-running Python processes that preserve state. Actors acquire resources for their lifetime. Thus, if multiple stateful UDFs require the same resource, multiplexing is possible by sharing an actor pool between UDFs.
For each physical operator, the Ray Data scheduler keeps a queue of ready partitions, either a Dataset.read input, e.g., a batch of input filenames, or an intermediate partition. We use Ray as a decentralized dataplane, so that the Ray Data scheduler only needs to keep references to partitions, not the physical data [44]. In a loop, the scheduler scans the current resource availability, chooses a ready partition, then passes it by reference to a new Ray task along with a description of the physical operator to execute. We describe the policy in Section 4.3.
We build upon and extend Ray’s fault tolerance. Ray provides lineage-based recovery and stores objects in separate processes from the workers, so that individual executor failures do not impact materialized partitions [44]. If a materialized partition is lost due to node failure, Ray automatically recreates it by resubmitting the tasks in its lineage. In Section 4.2.1, we describe the extensions we made to Ray’s recovery method to support streaming repartition.
@@ 4.1 Query Planning
The query planner takes as input a logical DAG and parameters such as the target maximum partition size. It generates a physical DAG, then applies a series of optimizations on the physical DAG, for example, adjacent operators that require the same resources are fused into one physical operator.
During execution, tasks may dynamically repartition the data. The initial number of partitions to use for a read operator is decided by the query planner. It needs to be sufficiently large to utilize all available execution slots (usually CPUs), but not so large that each partition is tiny, which can increase system overheads. It is also upper-bounded by the number of input files. By default, we aim to produce partitions that are 1—128 MB in size. We compute the initial number of partitions based on the following heuristics: the number of initial execution slots, the estimated output size of the read operator, and the user-requested value, if any.
Some Dataset consumption APIs (Table 2) induce transformations on the physical DAG. The Dataset.write call is appended to the DAG as a map that writes items to external storage. The iter API returns a stream of output records. Under the hood, this is implemented by fetching and buffering output partitions. The iter_split call shards the outputs into N streams, each of which can be passed to a different process. This is useful for cases such as distributed data-parallel training where the dataset is sharded among N trainers. To implement iter_split, the query planner launches a Ray actor before execution to coordinate the dynamic partition assignment. The stream readers fetch and buffer output partitions from this coordinator actor. Partitions are passed by reference to avoid coordinator bottleneck.
@@ 4.2 Query Execution
During execution, the Ray Data centralized scheduler has a global view of the executing tasks and the available resources. The scheduler repeatedly executes the following loop:
  • Wait for an executing task to materialize an output partition. Tasks may produce multiple output partitions; if this was the last, then mark the task’s resources as free.
  • While there are free resources and ready partitions, launch new tasks using the policy described in Section 4.3. Mark the task’s required resources as used.
The scheduler passes in the task description: (1) a closure of the physical operator to execute, (2) references to the task’s input partition(s), and (3) the target output partition size. We leverage Ray to ensure that the task’s input partition(s) are made local to its executor.
4.2.1 Streaming repartition
Although the query planner makes an initial estimate of the number of output partitions to use, this value may not be optimal. When a physical operator’s outputs are much larger or smaller than its inputs, the initial partitioning will produce too-large or too-small inputs. Too-large partitions can cause out-of-memory failures from buffering many records at the executor. On the other hand, too-small partitions are also inefficient; while the Ray Data scheduler only stores references to partitions, it still must manage some metadata per partition and carry out RPCs to schedule corresponding task(s).
Ray Data introduces a technique called streaming repartition to handle such cases (Figure 3(b)). To support this, we extend Ray with remote generator tasks, which enable Ray tasks to produce a dynamic number of outputs, and to pipeline execution with the task’s caller (the Ray Data scheduler). Whenever the task produces a new output, it notifies the Ray Data scheduler via RPC. Upon receipt, the scheduler can immediately launch a downstream task. Meanwhile, the upstream task can continue producing its next output partition.
Ray Data tasks take a target partition size from the scheduler and determine locally how to partition their outputs. When executing a task, the worker accumulates processed rows to a local output buffer. Once the output buffer exceeds the maximum target partition size (128 MB by default), it uses the yield keyword in Python to serialize and materialize the partition in Ray’s object store. If a logical operator produces much less data than it consumes, a task may produce a too-small partition. To handle this, we simply coalesce partitions by having the Ray Data scheduler pass multiple partitions from different upstream tasks to a single downstream task.
@@ 4.2.2 Failure recovery
Ray provides automatic recovery for objects (intermediate partitions) as long as (1) the driver is alive, (2) the tasks that created them are deterministic and side effect-free [44], and (3) the task arguments and outputs are immutable. For generator tasks, (3) is no longer true, because we do not know at submission time how many outputs the task will produce. However, we note that if (2) is true, then streaming repartition can be made deterministic. In particular, given a target partition size, a pure transform, we ensure that a Ray Data task will produce the same stream of output partitions if executed on the same input partition(s).
To support failure recovery for generator tasks, we modify Ray’s recovery subsystem to handle tasks with an unknown number of outputs. Generator tasks are initially launched with an unknown number of outputs. On the first successful execution, the task’s caller records the number of outputs that the task produces. If any of the task’s outputs are lost, we recover by re-executing the entire task. If the task produces a different number of outputs, we throw an error.
Similar to other batch and stream processing systems [9, 34, 47, 30], if the centralized scheduler dies, Ray garbage-collects the job and it must be re-executed from the beginning. In the future, this case can be optimized through known techniques for asynchronous global checkpointing [2, 9].
4.3 Adaptive Scheduler
The goal of the scheduler is to minimize the job completion time while keeping the total memory usage of intermediate data below the system limit. To achieve this, the scheduler uses a principle of equalizing the processing rates of each operator (in bytes per second). Intuitively, if processing rates are not equal, slower operators will accumulate pending inputs, eventually exhausting the memory buffer. Each operator’s processing rate can be controlled by deciding how many tasks to run in parallel for that operator. The processing rates are estimated online using run-time statistics, including operator task durations and average task input–output data size ratios, because these properties are difficult to predict ahead of time, and could vary depending on the actual data being processed.
4.3.1 Algorithm
The input to the scheduler is as follows:
  • The physical DAG of operators (§ 4.1). Each operator can process data items in parallel tasks and is annotated with its resource requirements, e.g., {GPU:1}.
  • Resource limits: the total number of CPU, GPU, or custom resource slots, and $totalMemoryCapacity$ of the system.
@@ 4.3.1 Algorithm
4:     Update budget $\triangleright$ Algorithm 2
@@
6:         Launch task of $source$
@@
9:     $Q\leftarrow\emptyset$triangleright Set of qualified operators
10:     for each operator $op$ in DAG do
11:         if hasInputData($op$) and
12:     hasAvailableResources($op$) and
13:     hasOutputBufferSpace($op$then
14:              $Q\leftarrow Q\cup\{op\}$
15:         end if
16:     end for
17:     if $Q\neq\emptyset$ then
18:         $selected\leftarrow\operatorname*{arg\,min}_{op\in Q}\text{bufferedOutputsSize}(op)$
19:         Launch task of $selected$
20:     end if
21:end while
Algorithm 1 describes the scheduling loop in step (3) of Figure 5. From all qualifying operators, it picks the one with the least amount of data accumulated in its output buffer. The intuition is that this operator would be producing data at a slower rate than it is consumed, and thus needs more parallelism. This policy works well for equalizing the operator processing rates when there is enough memory to store the intermediate data while all executors are utilized.
However, when memory is constrained, the policy (without lines 4–8) will result in resource under-utilization, as seen in Figure 4(a). The ideal solution is to keep the pipeline full and start source tasks, i.e. tasks for the source operator, as early as possible, as in Figure 4(b). To achieve this, lines 4–8 add a higher-priority optimistic policy for scheduling source tasks, described further in the next section.
4.3.2 Input Rate Control
The input rate, i.e. the rate at which source tasks are scheduled, must approximate the pipeline’s overall throughput: If the source tasks are launched too slowly, the downstream operators will have no input to process, wasting compute resources. Conversely, if the source tasks are launched too aggressively, then these tasks may starve downstream operators and eventually cause slowdowns from back-pressuring of the source operator and/or spilling to disk.
We use a dynamic memory budget algorithm to regulate the rate at which source tasks [^1] For DAGs with multiple sources, the launch rate for each operator should be proportional to their data output size. are launched. Intuitively, the budget is an optimistic estimate of the memory available for new data partitions to enter the system. When a source task is launched, we deduct its estimated output size from the budget. At every second, the budget is updated using Algorithm 2, which estimates the rate at which data leaves the pipeline.

### v5

@@ 3 Overview: The Streaming Batch Model
Streaming batch execution (Figures˜2(c) and 3c) presents two challenges: (1) supporting dynamically sized partitions, and (2) dynamically assigning cluster resources to partitions.
Challenge 1: Dynamic partitioning.
The throughput and memory footprint of a batch system is sensitive to the partition size. Our goal is to achieve memory-efficient pipelining without requiring users to tune a static partition plan. In particular, we want to allow the user to set only a target partition size in bytes, and the system should automatically partition intermediate data records at run time according to their actual memory usage. This is challenging because then the number of partitions is not known until run time. Meanwhile, batch systems typically require logging the partition plan before execution to enable lineage-based recovery.
(a) Static repartition.
(b) Dynamic repartition.
Figure 4: Arrows represent data dependencies. (a) Batch systems allow repartitioning, but it must be specified by the user before execution. Execution across operators cannot be pipelined. (b) Streaming batch systems dynamically and automatically repartition outputs to reduce peak memory usage. This also allows the next operator to start in parallel with A1.
@@
Figure 5: Scheduling under memory pressure. Green represents CPU executors’ local memory capacity (1 partition per CPU, 2 total). Pink represents the system’s shared memory capacity for intermediate data (1 partition total). Arrows represent data dependencies. (1) indicates a period where executors must stall and buffer outputs locally until there is sufficient space in shared memory. (a) further stalls CPU0 during (2) until there is enough shared memory to schedule A3, while (b) schedules A3 as soon as possible.
[…]
To reduce load on the scheduler, we use Ray so that executors only need to return references to their output partitions back to the scheduler, not the data [41]. However, Ray requires tasks to return all outputs at once, and lineage-based recovery requires the number of outputs to be known ahead of time. We extend Ray’s task scheduling and lineage subsystems to allow a task to stream its returned references (§ 4.2.2).
Challenge 2: Dynamic resource assignment.
The centralized scheduler maintains a global view of the current partition plan (Figure˜2(c)) and the available resources, allowing it to quickly reallocate resources to different operators at partition boundaries and implement cluster-level scheduling policies such as leveraging data locality to reduce data movement for distributed data-parallel training. However, efficient resource allocation for a heterogeneous pipeline is challenging and requires joint control with memory allocation.
For example, consider Figure˜5(a). Each CPU executor has local memory capacity (green) for 1 partition, and the system has shared memory capacity (pink) for 1 partition. In phase (1), A2 stalls and buffers its outputs locally until B1 completes. A conservative scheduler also waits for phase (2) before scheduling A3, to avoid stalling CPU0. Meanwhile, Figure˜5(b) reduces overall run time by scheduling A3 optimistically so that it finishes simultaneously with B2. Applying such an optimization successfully would ideally require knowing A3’s run time and memory footprint.
Section˜4.3 introduces our adaptive scheduler, which provides both policies in Figure˜5. We use a conservative backpressure mechanism similar to stream processing systems to implement Figure˜5(a). For Figure˜5(b), we use profile-guided optimization to estimate future memory availability, enabling robust optimistic scheduling. Because we use a centralized scheduler, both policies are simple to specify.
3.1 The Dataset API
A Dataset represents an application pipeline. Datasets are lazily created, by reading files or applying transforms to an existing Dataset (Table˜2). A Dataset is materialized through a write to storage, by iterating over the items in memory, or via cache.
A key part of the API is the ability to express resource requirements. Resource requirements are a map from resource name to float value and may be passed as an option to each transform. By default, each transform requires 1 CPU. Resource names can be CPU, GPU, or a custom resource label.
@@ 3.1 The Dataset API
cache Materialize all items and cache in memory.
@@
The map transforms take a stateless user-defined function (UDF) as an argument. For operations that require significant initialization time, such as a model loaded into GPU memory, we also support UDFs with read-only state. Like other lineage-based systems, we assume that UDFs are pure.
Figure 6: Ray Data architecture overview. Ray Data executes as a Ray library. The Ray Data scheduler maintains the partition metadata (Figure˜2(c)) and dispatches tasks to Ray workers (dashed arrows).
@@ 3.2 Executing a Ray Data Program
After creating a Dataset, the user triggers execution by calling one of the consumption APIs, as seen in (1) in Figure˜6. The Dataset is represented as a DAG of logical operators (Figure˜1). The system’s query planner then compiles this logical DAG into a DAG of physical operators ((2) in Figure˜6). The query planner applies operator fusion and decides the number of partitions to use for the first operator (§ 4.1).
Ray Data uses Ray as a task backend, storing intermediate data partitions in Ray’s distributed object store. During execution, the Ray Data scheduler maintains a global view of the current partition metadata (Figure˜2(c)), including a reference to the remote data. The scheduler then dispatches each physical operator as a series of Ray tasks ((3) in Figure˜6). Stateless UDF tasks are scheduled directly by Ray and can run on any Ray worker. For each stateful UDF, Ray Data creates a pool of Ray “actors”: stateful workers that acquire resources for their lifetime. If multiple stateful UDFs require the same resource, multiplexing is possible by sharing an actor pool between UDFs. Note that while actors hold read-only application state such as a GPU model, they do not store any system state, so any stateful UDF task can execute on any actor in the pool. This enables simple load balancing across workers. Furthermore, reallocating an operator’s resources and scaling the cluster up or down simply requires updating the Ray Data scheduler’s map of available resources.
The Ray Data scheduler runs a continuous loop alternating between waiting for a dispatched task to materialize an output partition and release its resources, and choosing a new task to run on any available resources. In principle, one could instead rely entirely on Ray’s scheduler by submitting the entire task graph in Figure˜2(a) upfront. However, this would prevent dynamic partitioning and naive application of Ray’s scheduler results in poor performance (Section˜5.3.3), as Ray cannot easily implement the application-aware policies that we present in Section˜4.3.
@@ 4.1 Query Planning
The query planner transforms a logical DAG (Figure˜1) into a physical DAG (Figure˜2(a)), then applies optimizations such as operator fusion and the initial partitioning for read. We choose enough partitions to use all available execution slots (usually CPUs), but not so many that each partition is tiny, which can increase system overheads. If the estimated output size is known for a particular file type, then we also aim to produce partitions that are 1—128 MB in size (Section˜5.3.2). Note that the system is robust to the initial number of partitions because we also dynamically repartition at run time.
Dataset consumption APIs (Table˜2) are also incorporated into the physical DAG. The Dataset.write call is implemented with map. The iter API returns a stream of output records. This is implemented by fetching and buffering materialized output partitions. The iter_split call shards the outputs into N streams, each of which can be passed to a different reader process. This is useful for cases such as distributed data-parallel training where the dataset is sharded among N trainers. To implement iter_split, the query planner uses an additional Ray actor to coordinate dynamic assignment of materialized output partitions to readers.
@@ 4.2 Query Execution
During execution, the Ray Data scheduler maintains a global view of materialized partitions (Figure˜2(c)), executing tasks and available resources. For each partition, the Ray Data scheduler maintains the following metadata: number of rows, size in bytes, and node location (used for data locality). For each operator, the scheduler maintains a queue of input partitions, stored as Ray references. It then repeatedly executes this loop:
  • Wait for an executing task to materialize an output partition. Push the partition reference onto the downstream operator’s input queue. Tasks may produce multiple output partitions; if this was the last, free the task’s resources.
  • While there are free resources and queued input partitions, select a physical operator to run using the policy described in Section˜4.3. Launch a new task for that operator. Mark the task’s required resources used.
To launch a new task, the scheduler passes a closure of the physical operator to execute and references to the task’s input partition(s). Ray ensures that all references are replaced by the physical data on the task executor. Typically, the scheduler will delete a reference as soon as its downstream task is submitted, signaling to Ray that the data can be garbage-collected after task execution. To implement Dataset.cache(), the scheduler simply keeps a reference to all output partitions of the corresponding operator.
4.2.1 Dynamic repartitioning
The query planner’s initial estimate of the number of partitions to use may not be optimal. Ray Data uses dynamic repartitioning to handle such cases. To support this, we extend Ray with generator tasks, enabling remote tasks to produce a dynamic number of outputs, and to pipeline execution with the task’s caller (the Ray Data scheduler). Whenever the task yields a new output partition, it notifies the Ray Data scheduler via RPC. Upon receipt, the scheduler reruns the scheduling loop and can launch a downstream task. Meanwhile, the upstream task continues producing its next output.
Ray Data executors take a target partition size from the scheduler and determine locally how to partition task outputs. A Ray Data task is an iterator that applies the given transform over its input partition(s). The executor yields from this iterator until the accumulated buffer exceeds the maximum target partition size (128 MB by default). Then, the Ray Data task flushes its buffer as a partition in Ray’s object store. For example, for a video decoding transform, Ray Data could assign one video per task, and each task would yield a stream of video frames. The Ray Data executor would then dynamically batch the frames into a stream of output partitions. Once the task finishes, all corresponding system state is removed from the executor.
Operators that produce much less data than they consume can produce too-small partitions. To handle this, the Ray Data scheduler coalesces partitions by passing multiple partitions from the upstream operator to a single task, up to the maximum target partition size.
For ML workloads, users often want control over GPU batch size. Thus, we expose a map_batches API that accepts a batch transform function and a target batch size. Ray Data tasks ensure the batch size by iterating over slices of the input partition(s). The Ray Data scheduler also coalesces too-small partitions to the desired batch size.
@@ 4.2.2 Failure recovery
Ray provides automatic recovery for objects (intermediate partitions) as long as (1) the driver is alive, (2) the tasks that created them are deterministic and side effect-free [41], and (3) the task arguments and outputs are immutable. For generator tasks, (3) is no longer true, because we do not know at submission time how many outputs the task will produce. However, we note that if (2) is true, then dynamic repartitioning can be made deterministic. In particular, given a target partition size and a pure transform, we ensure that a Ray Data task will produce the same stream of output partitions if executed on the same input partition(s).
To support failure recovery for generator tasks, we modify Ray’s recovery subsystem to handle tasks with an unknown number of outputs. On the first successful execution of a generator task, the task’s caller records the number of outputs that the task produces. If any of the task’s outputs are lost, we recover by re-executing the entire task. If the task produces a different number of outputs, we throw an error.
Similar to other batch and stream processing systems [7, 32, 43, 28], if the centralized scheduler dies, Ray garbage-collects the job and so Ray Data must re-execute from the beginning. Rollback can be reduced via checkpointing.
4.3 Scheduler Policy
The scheduler policy takes as input:
  • The physical DAG of operators (Figure˜6), each annotated with a resource requirement, e.g., {GPU:1}.
  • Cluster resources: for each node, the number of CPU, GPU, or custom resource slots, and the memory capacity of the shared memory pool that stores partitions.
The scheduler policy decides the current mapping between cluster resources and operator tasks/partitions. Each time a task completes, the policy can allocate the freed resources to a new operator task. The policy can also choose to allow more than one outstanding task per cluster resource. This is beneficial for cases such as large language model (LLM) inference, which batches continuously across sequences for better throughput [42]. With multiple tasks in flight to an LLM replica, batching can be applied to sequences across partitions in addition to within a partition (Section˜5.1.1).
@@ 4.3 Scheduler Policy
4:  Update budget $\triangleright$ Algorithm˜2
@@
6:   Launch task of sourcesource
@@
9:  $Q\leftarrow\emptyset$ $\triangleright$ Set of qualified operators
10:  for each operator opop in DAG do
11:   if hasInputData(opop) and   hasAvailableResources(opop) and   hasOutputBufferSpace(opop) then
12:    $Q\leftarrow Q\cup\{op\}$
13:   end if
14:  end for
15:  if $Q\neq\emptyset$ then
16:   $selected\leftarrow\operatorname*{arg\,min}_{op\in Q}\text{bufferedOutputsSize}(op)$
17:   Launch task of selectedselected
18:  end if
19:end while
The scheduler enforces the shared memory capacity as a hard limit, using either policy in Figure˜5. We implement the pessimistic policy (Figure˜5(a)) by prioritizing operators with shorter output queues and stalling in-progress tasks when the global memory limit has been reached. This is similar to the backpressure mechanism used in streaming systems. Next, we describe the optimistic policy that minimizes stalling based on run-time profiling.
4.3.1 Optimistic policy
The optimistic policy (Figure˜5(b)) aims to minimize stalling and therefore job completion time by estimating when memory will become available in the future. To achieve this, the scheduler estimates and then equalizes the processing rates of each operator (in bytes per second). Intuitively, if processing rates are not equal, slower operators will accumulate pending inputs, eventually exhausting memory. The optimistic policy estimates operator processing rates online using run-time statistics: task durations and ratio of input:output size. The policy controls operator processing rates by deciding how many tasks to run in parallel for that operator.
Algorithm˜1 describes the scheduling loop for the optimistic policy. Like the pessimistic policy, the optimistic policy prioritizes the operator with the least amount of buffered outputs. However, when memory is constrained, the policy (without lines 4–8) will result in resource under-utilization, as seen in Figure˜5(a). The ideal solution is to keep the pipeline full and start source tasks, i.e. tasks for the first operator, as early as possible, as in Figure˜5(b). To achieve this, lines 4–8 add a higher-priority optimistic policy for scheduling source tasks, described next.